# Data-Driven High-Dimensional Statistical Inference with Generative Models

### v1

@@ 4.6 Results
We additionally implement the likelihood envelope procedure described in Section 3.1, repeating the fit with the different bootstrapped models. The envelope is seen to increase the signal strength uncertainty by approximately $\sim 40\%$. An example for the fit results on a single mock dataset is shown in Fig. 8. It is unclear how much of the variation from the different bootstrapped models is coming from the finite statistics in the SB regions, and how much from intrinsic stochasticity of the generative model training process. If training hyperparameters were optimized, it is possible that stability could be improved, leading to a significant reduction in this uncertainty.
@@
Figure 8: A demonstration of our ’envelope of likelihoods’ approach on a single mock dataset. The final HI-SIGMA result is taken as the envelope of the different zero-aligned $\Delta\ln(\mathcal{L})$ of the five models trained on different bootstrapped background samples.
Full coverage studies with large samples of independent datasets would be necessary to assess whether the envelope uncertainty is appropriate, but we performed some basic checks based on the 5 mock datasets we have available. Of the 25 fit results from each of the separate bootstrapped models, in 16/25 (64%) the true signal strength is within the $1\sigma$ confidence interval (CI). The fit results were roughly evenly split between being above and below the true value, with 14/25 (56%) fits being above and the rest below. These results are suggestive there is no significant bias or undercoverage for the individual model fits. With the envelope of likelihoods approach, 4/5 fits contain the true value within the $1\sigma$-CI, with the remaining fit having a pull of $1.2\sigma$. 2/5 fits have a best fit signal strength above the true value. Further studies would be necessary to determine whether the envelope approach leads to appropriate coverage.
@@
The resulting impact of the shape systematic inclusion is shown in Fig. 9. In practice, we observe that the introduction of the shape uncertainty is noticeable but minor and affects the location of the best fit signal strength more than the associated uncertainty. Thus, to fully showcase its impact we combine it with the model uncertainty captured by bootstrapping. We observe how the introduction of the shape systematics increases the total uncertainty as expected.
@@ 5.1 Optimality
As demonstrated in Section 4.6, if using the same input variables, the HI-SIGMA approach offers a significant sensitivity gain as compared to the cut and fit approach and a moderate sensitivity gain as compared to fits to a binned classifier score. Additionally, because the classifier based approaches are typically trained fully on MC, data-MC differences could cause degradation of the sensitivity. Because classifiers are generally trained solely on simulation, they learn the optimal observable to distinguish simulated background and signal. However if the true background differs from the simulation, the classifier score observable will no longer be optimal for distinguishing true background and signal. In the HI-SIGMA approach, the background probability density is learned from the data itself, making it unreliant on data-MC mismodelings of the background. Furthermore, in the HI-SIGMA approach, all nuisances are profiled over within the full multidimensional space, allowing them to be optimally constrained by the data. This makes the HI-SIGMA approach maximally ’systematics-aware’, as compared to classifiers which generally lose their optimality in the presence of systematic uncertainties and much be trained in specific ways to mitigate this effect ATLAS:2024ynn ; CMS:2025cwy .
@@ 5.2 Data-driven Background Estimation
The HI-SIGMA approach naturally incorporates a data-driven background estimate, and our initial studies demonstrate how its uncertainties can be parameterized. In the cut+fit approach, the background estimate is typically straightforward, just done through a parametric mass fit. However caution must be taken to ensure the classifier score cut does not sculpt the mass distribution and cause a bias. This can be difficult to verify for very tight classifier score cuts, often used in high-purity signal regions, in which there are few remaining events. In the binned classifier score approach, the background estimate is less straightforward, generally some analysis-specific methodology must be employed and may not be possible in all cases. Within a resonance search of the type considered in the HI-SIGMA approach, some sort of interpolation of the background distribution from the sideband regions is needed, which may only be possible for specific types of generative models. The current $bb\gamma\gamma$ analyses from ATLAS and CMS ATLAS_bbgg ; ATLAS_bbgg_run3 ; CMS_bbgg use a cut+fit strategy, likely because of these difficulties.
@@ 5.4 Comparison to SBI
The largest advantage for the classifier-based approach of SBI may be practicality. It is common wisdom that classifiers are easier to train than generative models. One explanation is that training a classifier on $N$ input features encodes a function $\mathbb{R}^{N}\to\mathbb{R}$, whereas typical generative models (such as normalizing flows or diffusion models) learn a more complex $\mathbb{R}^{N}\to\mathbb{R}^{N}$ function. It is therefore unsurprising that classifiers take less computational resources to train, can be trained with fewer examples, and can more easily scale to large $N$. Further research is be required to assess how large a feature set can be accurately modeled in the HI-SIGMA approach, and assess how many features are required for optimal inference in many analyses. However, given previous applications of generative models, its likely that HI-SIGMA could scale to $O(10)$ dimensional distributions which is sufficient for many analyses.
@@ 6 Conclusions and Outlook
In this paper we have introduced HI-SIGMA, a new analysis methodology tailored to resonances, enabling unbinned high dimensional statistical inference with data-driven background modeling. Using as an example di-Higgs measurement in the $bb\gamma\gamma$ final state, we have demonstrated that HI-SIGMA provides sensitivity gains over commonly employed classifier-based analysis strategies, while providing a data-driven background estimate and maintaining interpretability. Because HI-SIGMA performs statistical inference in a high-dimensional space, it should offer significant gains over traditional methods in multi-parameter inference. This capability could be crucial for LHC analyses seeking to constrain multiple effective field theory operators, as was proposed for classifier-based SBI methods Brehmer:2018eca ; Brehmer:2018kdj . Many of the flagship LHC measurements in coming years will feature the narrow resonances, such at the Higgs, and require data-driven backgrounds, and are thus amenable to the HI-SIGMA.
A key question for future work would be to assess how high-dimensional feature spaces can the HI-SIGMA approach can still offer accurate density estimates. For scaling to very large dimensionalities it may be interesting to explore diffusion models, which have been shown to achieve high dimensional density estimates on $O(100)$ dimensional particle physics data Mikuni:2023tok . If the HI-SIGMA approach is seen to not be able to accurately model very large dimensionalities, it could be used in conjunction with a classifier in a hybrid approach. A selection using a classifier trained on the whole feature space could be applied, and then HI-SIGMA used to fit the most important features on the remaining events. Further work will also be needed to appropriately assess the uncertainties on this method of background estimation.

### v2

@@ 5.2 Performance Quantification
5.3 Robustness to Simulation Mismodeling
Another advantage of HI-SIGMA as compared to classifiers is that it learns the background density directly from data and thus avoids any loss of optimality due to imperfect MC background simulation. Since the classifiers are trained on MC, any data-MC difference will result in a suboptimal classifier score and a consequent degradation in sensitivity. This effect persists even if data-driven methods are employed to estimate the background for the classifier-based analyses. To illustrate this, we perform a mock study where we construct a training sample for our classifier with slightly distorted features and then use the resulting trained classifier for inference on un-distorted event samples.
Supervised classifiers are trained using these samples using the same architecture and hyperparameters discussed in Section 4.5. These classifiers are used to perform inference on the undistorted event samples. The results of these inferences are shown in Fig. 10. We observe that the training distortion degrades the sensitivity of the classifier-based methods by $\sim$20%. In contrast, as HI-SIGMA does not rely on background simulation at all, it is not affected by these distortions.
For the cut+fit approach, because of our limited dataset size and reduced feature space, the optimal classifier selection ends up being quite loose, rejecting only 95% of background events. The efficiency of this loose selection is therefore relatively unaffected by the MC distortion, even though the classifier itself has reduced performance. For the comparison in Fig. 10 we therefore employ a slightly tighter classifier cut, rejecting 99% of background, to demonstrate the effect. This tighter selection has slightly better $s/\sqrt{b}$ but slightly worse sensitivity ($\pm 0.15$) as compared to the optimal selection ($\pm 0.14$), due to larger uncertainties on the background shape arising from the reduced event count. We note that current $bb\gamma\gamma$ analyses from ATLAS and CMS ATLAS_bbgg; ATLAS_bbgg_run3; CMS_bbgg cut very tightly on the classifier score, leaving only ${\sim}10$ events in the final signal region, and therefore could certainly be impacted by data-MC differences.
Figure 10: A comparison of the precision of the classifier-based approaches trained with distorted MC as compared to their original performance and HI-SIGMA. The curve shown for the DNN score fit is based on the 50 bin version. The mismodelings of the background MC used in the classifier training reduces the sensitivity of the DNN-based approaches by $\sim$}20%, whereas HI-SIGMA is unaffected.
5.4 Impact of Systematic Uncertainties
@@
We implement the likelihood envelope procedure described in Section 3.1, repeating the fit with the different bootstrapped models. The envelope is seen to increase the signal strength uncertainty by approximately ${\sim}40\%$. An example for the fit results on a single mock dataset is shown in Fig. 11. It is unclear how much of the variation from the different bootstrapped models is coming from the finite statistics in the SB regions, and how much from intrinsic stochasticity of the generative model training process. If training hyperparameters were optimized, it is possible that stability could be improved, leading to a significant reduction in this uncertainty.
@@
Figure 11: A demonstration of our ‘envelope of likelihoods’ approach on a single mock dataset. The final HI-SIGMA result is taken as the envelope of the different zero-aligned $\Delta\ln(\mathcal{L})$ of the five models trained on different bootstrapped background samples.
@@
We also implement the toy shape systematic described in Section 3.2. The resulting impact from its inclusion is shown in Fig. 12. In practice, we observe that the introduction of the shape uncertainty is noticeable but minor and affects the location of the best fit signal strength more than the associated uncertainty. Thus, to fully showcase its impact we combine it with the model uncertainty captured by bootstrapping. We observe how the introduction of the shape systematics increases the total uncertainty as expected.
@@
5.5 Coverage Studies
Full coverage studies with large samples of independent datasets would be necessary to assess whether the envelope uncertainty is appropriate, but we performed some basic checks based on the 5 mock datasets of size 40k.
First, we consider five different signal strengths, varying from zero injected signal events up to the 200 events considered in the previous studies. For each signal strength we inject a random selection of signal events into each mock dataset and perform inference. For simplicity, we include only the uncertainty associated with the finite training statistics, and do not consider the toy shape systematics. The results are shown in Figure 13.
Figure 13: The extracted number of signal events for different injection tests of HI-SIGMA on five independent datasets of size 40k. For larger signal injections, the measured values appear to adequately contain the true values, while for smaller signals a bias becomes more evident.
We see that for the larger signal injections of 100 and 200 events, HI-SIGMA seems to be performing well, with the true signal being within the uncertainties of most of the inferred signals and showing no obvious bias. However at low signal injections a bias begins to appear, where the inferred signals are consistently higher than the true signal values. In particular, for zero injected signal the bias appears to be larger than $1\sigma$.
It is likely that the bias arises from some mismodeling of the background by the normalizing flow model. If some region of phase space has an artificially high signal-to-background density ratio due to artificial background suppression, the fit would prefer to increase the fitted number of signal events in order to increase the likelihood of all background events that fall in that region. To check if this bias originates from issues in the interpolation of the background estimate from the sidebands into the mass signal region, separate background models were trained using events from the signal region. Inference performed with these background models were found to be similarly biased, indicating the bias does not appear related to the interpolation but instead seems to be a reflection of a limitation of the current density modeling. Further work would be needed to improve the background model and eliminate this bias.
To further test the coverage properties of HI-SIGMA at low signal strengths we perform a more thorough coverage test using small event samples. We split our available testing set of 200k events into 50 independent samples of 4k events. For each sample we inject a known number of signal events, perform inference with HI-SIGMA, and compare the measured signal strength to the true value. We repeat this procedure for two signal strengths: a background-only test, where no signal events are injected, and for an injection of 25 signal events. The latter injection corresponds to approximately $2.5\sigma$ of evidence, due to the reduced background statistics. For each measurement we compute the pull, defined as
$t=\frac{s_{\mathrm{meas}}-s_{\mathrm{true}}}{\sigma},$ (14)
for measured signal events $s_{\mathrm{meas}}$, true signal events $s_{\mathrm{true}}$ and measurement uncertainty $\sigma$. If HI-SIGMA yields an unbiased and properly calibrated estimator, the pull distribution will be a Gaussian with mean zero and standard deviation of one. The 1-$\sigma$ coverage fraction can be assessed as the fraction of measurements with $|t|<1$. A well-calibrated pull should yield a 1-$\sigma$ coverage fraction consistent with 68.3%.
In Table 1 we report these quantitative metrics for HI-SIGMA both with and without the finite training uncertainty described in Section 3.1. We observe that for all four cases the pull distribution indicates no significant bias. Proper coverage is achieved when including the finite training uncertainty.
The fact that no significant bias for the no signal case is observed on this smaller sample, in contrast to the apparent bias seen on the larger sample, further reinforces the hypothesis that the source of previously observed bias resides in the background mismodelling. The effect of this mismodeling becomes subdominant with respect to the increased statistical uncertainties of the smaller sample, ameliorating the bias. This is encouraging, as it implies that the bias in the background estimation is likely small, and that HI-SIGMA can perform reliable inference for moderate sample sizes of $\mathcal{O}(5\times 10^{3})$, which are realistic for many analyses.
We expect the particular range of trustworthy sample sizes to depend on the amount of training data and quality of training. In general, further improvements to the generative model used for the background estimate or better uncertainty parameterizations would allow unbiased inference on larger samples. For a realistic applications, such as a $bb\gamma\gamma$ analysis, a classifier selection could be applied prior to HI-SIGMA, and the threshold of the classifier chosen to result in a number of events amenable to accurate inference from HI-SIGMA. If the classifier selection does not sculpt the mass distribution or introduce sharp features that are difficult to capture in the background model, HI-SIGMA could be used in this signal-enriched dataset without any modifications.
Method Injection size Mean pull Std. dev. pull 1-$\sigma$ coverage frac.
HI-SIGMA (no sys) 0 $0.12\pm 0.25$ $1.72\pm 0.17$ $38\pm 9$%
25 $0.24\pm 0.15$ $1.03\pm 0.10$ $64\pm 11$%
HI-SIGMA (w/ sys) 0 $0.14\pm 0.15$ $1.05\pm 0.11$ $64\pm 11$%
25 $0.19\pm 0.12$ $0.87\pm 0.09$ $72\pm 12$%
Table 1: Coverage test performed on 50 independent datasets of 4k events each. The pull distribution mean, standard deviation and 1-$\sigma$ coverage fraction are reported for two different signal injections. The evaluations are performed both and without the systematic uncertainty arising from finite training statistics shown in Sec. 3.1. Uncertainties on the reported values originate from the limited number of toys. Good coverage properties are seen for both injection sizes with the inclusion of systematic uncertainties.
@@ 6.1 Optimality
As demonstrated in Section 5, if using the same input variables, the HI-SIGMA approach offers a significant sensitivity gain as compared to the cut and fit approach and a moderate sensitivity gain as compared to fits to a classifier score fit with a low number of bins. Additionally, because the classifier based approaches are typically trained fully on MC, data-MC differences could cause degradation of the sensitivity. In the HI-SIGMA approach, the background probability density is learned from the data itself, making it unreliant on data-MC mismodelings of the background. This was demonstrated in Section 5.3.
Furthermore, in the HI-SIGMA approach, all nuisances are profiled over within the full multi-dimensional space, allowing them to be optimally constrained by the data. This makes the HI-SIGMA approach maximally ‘systematics-aware’, as compared to classifiers which generally lose their optimality in the presence of systematic uncertainties and must be trained in specific ways to mitigate this effect ATLAS_SBI_methods; CMS:2025cwy.
@@ 6.2 Data-driven Background Estimation
The HI-SIGMA approach naturally incorporates a data-driven background estimate, and our initial studies demonstrate how its uncertainties can be parameterized. In the cut+fit approach, the background estimate is typically straightforward, just done through a parametric mass fit. However caution must be taken to ensure the classifier score cut does not sculpt the mass distribution and cause a bias. This can be difficult to verify for very tight classifier score cuts, often used in high-purity signal regions, in which there are few remaining events. In the binned classifier score approach, the background estimate is less straightforward. Generally some analysis-specific methodology must be employed and a reliable strategy may not be possible in all cases. Within a resonance search of the type considered in the HI-SIGMA, some sort of interpolation of the background distribution from the sideband regions is needed, which may only be possible for specific types of generative models. The current $bb\gamma\gamma$ analyses from ATLAS and CMS ATLAS_bbgg; ATLAS_bbgg_run3; CMS_bbgg use a cut+fit strategy, despite the known sensitivity increase from binned classifier score fits, likely because of these difficulties.
@@ 6.4 Comparison to SBI
The largest advantage for the classifier-based approach of SBI may be practicality. It is common wisdom that classifiers are easier to train than generative models. One explanation is that training a classifier on $N$ input features encodes a function $\mathbb{R}^{N}\to\mathbb{R}$, whereas typical generative models (such as normalizing flows or diffusion models) learn a more complex $\mathbb{R}^{N}\to\mathbb{R}^{N}$ function. It is therefore unsurprising that classifiers take less computational resources to train, can be trained with fewer examples, and can more easily scale to large $N$. Further research is required to assess how large a feature set can be accurately modeled in the HI-SIGMA approach, and assess how many features are required for optimal inference in many analyses. However, given previous applications of generative models, its likely that HI-SIGMA could scale to $O(10)$ dimensional distributions which is sufficient for many analyses.
@@ 7 Conclusions and Outlook
In this paper we have introduced HI-SIGMA, a new analysis methodology tailored to resonances, enabling unbinned high dimensional statistical inference with data-driven background modeling. Using as an example di-Higgs measurement in the $bb\gamma\gamma$ final state, we have demonstrated that HI-SIGMA provides sensitivity gains over commonly employed classifier-based analysis strategies, while providing a data-driven background estimate and maintaining interpretability. Because HI-SIGMA performs statistical inference in a high-dimensional space, it should offer significant gains over traditional methods in multi-parameter inference when the feature space dimensionality is $\mathcal{O}(10)$. This capability could be crucial for LHC analyses seeking to constrain multiple effective field theory operators, as was proposed for classifier-based SBI methods Brehmer:2018eca; Brehmer:2018kdj. Many of the flagship LHC measurements in coming years will feature the narrow resonances, such at the Higgs, and require data-driven backgrounds, and are thus amenable to the HI-SIGMA.
A key question for future work would be to assess the impact of the dimensionality of the feature space on the accuracy of the HI-SIGMA approach. Further validations of the accuracy of the density estimation, more rigorous uncertainty quantification methods on the density estimates, and full coverage studies would be needed prior to deployment. For scaling to very large dimensionalities it may be interesting to explore diffusion models, which have been shown to achieve high dimensional density estimates on $O(100)$ dimensional particle physics data Mikuni:2023tok. If the HI-SIGMA approach is seen to not be able to accurately model very large dimensionalities, it could be used in conjunction with a classifier in a hybrid approach. A selection using a classifier trained on the whole feature space could be applied, and then HI-SIGMA used to fit the most important features on the remaining events. Further work will also be needed to appropriately assess the uncertainties on this method of background estimation.